# When and How: Learning Identifiable Latent States for Nonstationary Time Series Forecasting

### v1

@@ 3.3 Identification of Nonstationary Latent Variables ${\mathbf{z}}^{e}_{t}$
  • A3.3.1 ((Smooth and Positive Density:)) The probability density function of latent variables is smooth and positive, i.e. $p({\mathbf{z}}_{t}^{e}|{\mathbf{e}}_{t})>0$ over $\mathcal{Z}_{t}$ and $\mathcal{E}_{t}$.
  • A3.3.2 (Conditional independent:) Conditioned on ${\mathbf{e}}_{t}$, each ${\mathbf{z}}^{e}_{t,i}$ is independent of any other ${\mathbf{z}}_{t,j}^{e}$ for $i,j\in\{1,\cdots,n_{e}\},i\neq j$, i.e., $\log p({\mathbf{z}}_{t}^{e}|{\mathbf{e}}_{t})=\sum_{i=1}^{n_{e}}\log p(z_{t,i}^{e}|{\mathbf{e}}_{t})$.
  • A3.3.3 (Linear independence): For any ${\mathbf{z}}_{t}^{e}\in\mathcal{Z}^{e}_{t}\subseteq\mathbb{R}^{n_{e}}$, there exist $n_{e}+1$ values of ${\mathbf{e}}$, i.e., ${\mathbf{e}}_{j}$ with $j=1,2,..n_{e}+1$, such that these $n_{e}$ vectors $\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})-\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{0})$ are linearly independent, where the vector $\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})$ is defined as follows:
    $\small\mathbf{w}({\mathbf{z}}^{e}_{t},{\mathbf{e}}_{j})\!=\!\left(\frac{\partial\log p(z_{t,1}^{e}|{\mathbf{e}}_{j})}{\partial z_{t,1}^{e}},\cdots,\frac{\partial\log p(z_{t,n_{e}}^{e}|{\mathbf{e}}_{j})}{\partial z_{t,n_{e}}^{e}}\right)$ (6)
@@
Then, by learning the data generation process, ${\mathbf{z}}^{e}_{t}$ are subspace identifiable.
The detailed proof can be found in the Appendix E.3, which is similar to that of Theorem 3.2, and it also shows that the nonstationary latent variables are subspace identification with only $n_{e}+1$ prior environments.
In summary, we can show that $p({\mathbf{x}})$ is identifiable with the help of the aforementioned theoretical results.
@@ 4 Identifiable Latent States Model
In this section, we introduce the implementation of the identifiable latent states model (IDEA) as shown in Figure 3, which is built on a sequential variational inference module with an autoregressive hidden Markov Module for latent environment estimation. Moreover, we devise modular prior networks to estimate the prior of stationary and nonstationary latent variables in evidence lower bound.
Figure 3: The framework of the Identifying Disentnagled Latent State model. The latent variable encoder is used to extract ${\mathbf{z}}^{s}_{1:t}$ and ${\mathbf{z}}^{e}_{1:t}$ from ${\mathbf{x}}_{1:t}$. The latent forecasting module is used to estimate ${\mathbf{z}}^{s}_{t+1:T}$ and ${\mathbf{z}}^{e}_{t+1:T}$ from ${\mathbf{z}}^{s}_{1:t}$ and ${\mathbf{z}}^{e}_{1:t}$, respectively. And the future forecasting module is used for future prediction $\{\hat{x}_{t+1},\cdots,\hat{x}_{T}\}$. The historical latent environments $\{\hat{{\mathbf{e}}}_{1},\cdots,\hat{{\mathbf{e}}}_{t}\}$ are generated by the environment estimation module, and the future latent environments $\{\hat{{\mathbf{e}}}_{t+1},\cdots,\hat{{\mathbf{e}}}_{T}\}$ are generated by the environment prediction module. The nonstationary prior and the stationary prior are used to estimate the prior distribution of stationary and nonstationary latent variables for KL divergence.
@@ 4.1 Sequential Variational Inference Module for Time Series Data Modeling
According to the data generation process in Figure 2, we first derive the evidence lower bound (ELBO) as shown in Equation (7). Please refer to Appendix D for more details on the derivation of the ELBO.
@@
$\small\begin{split}\mathcal{L}_{KLD}^{s}\!=&D_{KL}\!(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}))\\ +&\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\!\Big{[}\!D_{KL}\!(q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}))\Big{]}\\ \mathcal{L}_{KLD}^{e}\!=&D_{KL}\!(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}))\\ +&\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\!\Big{[}\!D_{KL}\!(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t}))\Big{]},\end{split}$ (9)
in which $q({\mathbf{z}}_{1:t}^{s}|{\mathbf{x}}_{1:t}),\,\,q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}_{1:t}^{s}),\,\,q({\mathbf{z}}_{1:t}^{e}|{\mathbf{x}}_{1:t})$ and $q({\mathbf{z}}_{t+1:T}^{e}|{\mathbf{z}}_{1:t}^{e})$ are used to approximate the distribution; Therefore, the aforementioned approximated functions, historical decoder, and the future forecasting module can be formalized as follows:
$\small\begin{split}{\mathbf{z}}^{e}_{1:t}&=\psi_{e}({\mathbf{x}}_{1:t};\theta_{\psi_{e}}),\quad\quad\,\,\,\,\quad{\mathbf{z}}^{s}_{1:t}=\psi_{s}({\mathbf{x}}_{1:t};\theta_{\psi_{s}}),\\ {\mathbf{z}}^{e}_{t+1:T}&=T_{e}({\mathbf{z}}^{e}_{1:t};\theta_{T_{e}}),\,\,\,\,\quad\quad{\mathbf{z}}^{s}_{t+1:T}=T_{s}({\mathbf{z}}^{s}_{1:t};\theta_{T_{s}}),\\ \hat{{\mathbf{x}}}_{1:t}&=F_{x}({\mathbf{z}}^{e}_{1:t},{\mathbf{z}}^{s}_{1:t};\theta_{x}),\quad\hat{{\mathbf{x}}}_{t+1:T}=F_{y}({\mathbf{z}}^{e}_{t+1:T},{\mathbf{z}}^{s}_{t+1:T};\theta_{y}),\end{split}$ (10)
@@
$p({\mathbf{z}}^{s}_{1:t}),\,\,p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}),\,\,p({\mathbf{z}}^{e}_{1:t})$ and $p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})$ are the estimated prior distribution of stationary and nonstationary latent variables, which will be introduced in the subsection 4.2. Note that the environment estimation module does not appear in the ELBO explicitly, which is a part of the $p({\mathbf{z}}^{e}_{1:t}),p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})$ and will be also introduced in subsection 4.2. Please refer to the Appendix H for more details of the implementation of the proposed IDEA model.
4.2 Modular Architectures with Stationary and Nonstationary Priors
Previous time-series modeling methods based on the causality-based data generation processes usually require autoregressive inference and Gaussian prior (Fabius & Van Amersfoort, 2014; Kim et al., 2019; Zhu et al., 2020). However, these prior distributions usually contain temporal information and follow an unknown distribution. Simply assuming the Gaussian distribution might result in suboptimal performance of disentanglement. To solve this problem, we employ the modular neural architecture to estimate the prior distribution of stationary and nonstationary latent variables.
Modular Architecture for the Stationary Prior Estimation. we first let $\{r_{i}^{s}\}$ be a set of learned inverse transition functions that take the estimated stationary latent variables and output the noise term, i.e., $\hat{\epsilon}_{t,i}^{s}=r_{i}^{s}(\hat{z}_{t,i}^{s},\hat{{\mathbf{z}}}^{s}_{t-1})$  [^1] We use the superscript symbol to denote estimated variables. and each $r_{i}^{s}$ is modeled with MLPs. Then we devise a transformation $\phi^{s}:=\{\hat{{\mathbf{z}}}^{s}_{t-1},\hat{{\mathbf{z}}}^{s}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}^{s}_{t-1},\hat{\epsilon}^{s}_{t}\}$, and its Jacobian is $\small{\mathbf{J}_{\phi^{s}}=\begin{pmatrix}\mathbb{I}&0\\ *&\text{diag}\left(\frac{\partial r^{s}_{i}}{\partial\hat{z}_{t-1,i}^{s}}\right)\end{pmatrix}}$, where *** denotes a matrix. By applying the change of variables formula, we have the following equation:

### v4

@@ 3.1 Latent Environment Identification
  • A1 (Smooth and Positive Density:) The probability density function of latent variables is smooth and positive, i.e., $p({\mathbf{z}}^{e}_{t}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})>0$ over $\mathcal{Z}_{t}^{e},\mathcal{Z}_{t-1}^{e}$ and $\mathcal{Z}_{t-2}^{e}$.
  • A2 (Sufficient Variability of Historical Information:) For any ${\mathbf{z}}_{t}^{e}\in\mathcal{Z}_{t}^{e}\subseteq\mathbb{R}^{n_{e}}$, $\bar{{\mathbf{v}}}_{t-1,1},\cdots,\bar{{\mathbf{v}}}_{t-1,n_{e}}$ as $n_{e}$ vector functions in $z_{t-2,1},\cdots,z_{t-2,l},\cdots,z_{t-2,n_{e}}$ are linear independent, where $\bar{{\mathbf{v}}}_{t-2,l}$ are formalized as follows:
    $\bar{{\mathbf{v}}}_{t-2,l}=\frac{\partial^{2}\log p({\mathbf{z}}_{t}^{e}|{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e})}{\partial z_{t,k}^{e}\partial z_{t-2,l}^{s}}$ (5)
  • A3 (Sufficient Variability of Environments:) There exist two values of ${\mathbf{u}}=\{{\mathbf{z}}_{t-1}^{e},{\mathbf{z}}_{t-2}^{e}\}$, i.e., ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$, s.t., for any set $\mathcal{A}_{{\mathbf{z}}_{t}}\subseteq\mathcal{Z}_{t}$ with non-zero probability measure and $\mathcal{A}_{{\mathbf{z}}_{t}}$ cannot be expressed as $B_{{\mathbf{z}}_{t}^{s}}\times\mathcal{Z}_{t}^{e}$, for any $B_{{\mathbf{z}}_{t}^{s}}\subset\mathcal{Z}_{t}^{s}$, we have:
[…]
@@
Then, by learning the data generation process, ${\mathbf{z}}_{t}^{e}$ and ${\mathbf{z}}_{t}^{s}$ are block-wise identifiable.
Proof Sketch. The proof can be found in Appendix E.2 First, we construct an invertible transformation $h$ between the ground-truth latent variables ${\mathbf{z}}_{t}$ and the estimated ones $\hat{{\mathbf{z}}}_{t}$. According to the data generation process in Figure 2, we find that ${\mathbf{z}}_{t}^{e}$ is dependent on ${\mathbf{z}}_{t-2}$ while ${\mathbf{z}}_{t}^{s}$ is independent of ${\mathbf{z}}_{t-2}$ given ${\mathbf{z}}_{t-1}$. Hence we can construct a full-rank linear system, where the only solution of $\frac{\partial{\mathbf{z}}^{e}_{t}}{\partial\hat{{\mathbf{z}}}_{t}^{s}}$ is zero. Because of the invertibility of the Jacobian of $h$ and the variability of historical information, both ${\mathbf{z}}_{t}^{e}$ and ${\mathbf{z}}_{t}^{s}$ are block-wise identifiable.
Based on Theorem 3.1, we can make sure that the estimated $\hat{{\mathbf{z}}}_{t}^{e}$ contain all the information of the truth ${\mathbf{z}}_{t}^{e}$ and do not contain the information of stationary latent variables ${\mathbf{z}}_{t}^{s}$. So we can consider ${\mathbf{z}}_{t}^{e}$ as observed variables and further leverage the results from (Allman et al.,, 2009), as shown in Lemma 3.2.
Lemma 3.2.
(Identifiability of the latent environment ${\mathbf{e}}_{t}.$ (Allman et al.,, 2009)) Suppose the observed data are generated following the data generation process in Figure 3 and Equation (1)-(3). Then we further make the following assumptions:
  • A4 (Prior Environment Number:) The number of latent environments, $E$, is known.
  • A5 (Full Rank:) The transition matrix $\mathbf{A}$ is full rank.
  • A6 (Linear Independence:) For $e=1,2,\cdots,E$, the probability measures $\mu_{e}=p({\mathbf{z}}_{t}^{e}|e_{t})$ are linearly independence and for any two different probability measures $\mu_{i},\mu_{j}$, their ratio $\frac{\mu_{i}}{\mu_{j}}$ are linearly independence.
Then, by modeling the observations ${\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t}$, the joint distribution of the corresponding latent environment variables $p({\mathbf{e}}_{1},\cdots,{\mathbf{e}}_{t})$ is identifiable up to label swapping of the hidden environment.
Proof Sketch. First, given any three consecutive observations ${\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3}$ with the corresponding latent environments ${\mathbf{e}}_{1},{\mathbf{e}}_{2},{\mathbf{e}}_{3}$, we derive the joint distribution of $p({\mathbf{x}}_{1},{\mathbf{x}}_{2},{\mathbf{x}}_{3})$ to the product of three independent measures w.r.t. $p({\mathbf{e}}_{2})$. Sequentially, by employing the extension of Kruskal’s theorem (Kruskal,, 1977, 1976), the latent environment variables can be detected with identification guarantees. The detailed proof of Lemma 3.2 is provided in Appendix E.1.
3.2 Component-wise Identification of Stationary and Nonstationary Latent Variables
Based on the aforementioned theoretical results, we prove that the stationary and nonstationary latent variables are component-wise identifiable with the help of nonlinear ICA.
Lemma 3.3.
(Component-wise Identification of the stationary latent variables ${\mathbf{z}}^{s}_{t}$ and nonstationary latent variables ${\mathbf{z}}_{t}^{e}$.(Yao et al.,, 2021)) Following the data generation process in Figure 2, smooth and density as well as the similar sufficient variability assumptions of Theorem 3.1, we further assume that the latent variables are conditionally independent, ${\mathbf{z}}^{s}_{t}$ is component-wise identifiable.
Proof Sketch The proof can be found in the Appendix E.2, which contains the identification of the stationary and nonstationary variables, respectively. The proof of both types of latent variables is similar. Specifically, we first construct an invertible transformation $h$ between the ground-truth latent variables and estimated ones. Then we employ the variance of different environments to construct a full-rank linear system, where the only solution is zero.
3.3 Comparison with Existing Methods
Table 1: Attributes of causal representation learning theories. A check denotes that a method has an attribute or can be applied to a setting, whereas a cross denotes the opposite.
Methods
Partitioned
Subspace
Time-Delayed
Causal Relations
No Extra Assumptions
on Transition
HMNLICA
NCTRL
CtrlNS
UDA
Although recent advances (Song et al.,, 2024, 2023; Hälvä and Hyvarinen,, 2020) also achieve the identification for temporal representation under unknown nonstationarity, our method works under less restrictive conditions and better reflects real-world scenarios, as summarized in Table 1.
First, compared to existing methods, the proposed UDA is better suited for nonstationary time series forecasting, as it allows for partitioned subspaces. In contrast, methods like Hälvä and Hyvarinen, (2020) cannot be applied to the data generation process, as shown in Figure 2, because the causal relationships induced by stationary latent variables disrupt the conditional independence among nonstationary latent variables. Moreover, our method allows for time-delayed causal relationships among latent variables, specifically reflected in stationary latent variables.
Most importantly, compared with other methods, our approach does not impose additional assumptions on transitions among latent variables. For instance, Hälvä and Hyvarinen, (2020) assume the absence of time-delayed causal relationships, and Song et al., (2023) constrains transitions to a nonlinear Gaussian family with unique indexing (Balsells-Rodas et al.,, 2023). Furthermore, Song et al., (2024) further assume a sparse latent transition. These assumptions may not be met in real-world scenarios.
3.4 Discussion of Assumptions
For a better understanding of our theoretical results, we further provide detailed explanations and implications of the assumptions of these theories, as well as how they relate to the real-world time series data.
Smooth, Positive, and Conditional Independent Density. This assumption is commonly used in existing identification results (Yao et al.,, 2022, 2021). In real-world scenarios, a smooth and positive density implies continuous changes in historical information, such as temperature fluctuations in weather data. To achieve this, collecting a large amount of data is essential for accurately learning transition probabilities. Moreover, the conditional independent assumption is also common in identifying temporal latent processes (Kong et al.,, 2022; Li et al., 2023a, ). Intuitively, it means there are no immediate relations among latent variables. To satisfy this assumption, we can sample data at high frequency to avoid instantaneous dependencies caused by subsampling.
Sufficient Variability. The implications of the sufficient variability of historical information or environment in Theorem 3.1 are similar. It is also common in (Yao et al.,, 2022; Kong et al.,, 2022), reflecting that the influence of each latent variable on the observations is independent. This assumption is also a standard requirement for identifying nonlinear ICA (Allman et al.,, 2009; Hyvarinen and Morioka,, 2016; Hälvä and Hyvarinen,, 2020; Lippe et al., 2022b, ), ensuring a unique solution to the system of equations. Although this assumption is untestable, it can be assessed based on prior knowledge specific to the application.
Prior Environment Number. The prior environment number assumption implies that we can take the number of environments as prior knowledge. For example, we can know the number of actions of drivers.
Full Rank and Linear Linear Independence. The full-rank assumption implies that the state transition matrix of latent environment is full-rank, meaning the transition probability between any two environments is nonzero. However, if the collected data is insufficient, it may not capture all environment transitions, causing this assumption to break down. To meet this assumption, we should collect as much data as possible across diverse environmental conditions.
Linear Independent. The Linear Independence assumption aligns closely with the concept of sufficient variability. It implies that when the environment changes, the resulting variations in observed variables are significant. For example, the effects of speeding and braking on a car’s speed are fundamentally different, reflecting distinct influences from the underlying environment.
@@ 4 Identifiable Latent States Model
In this section, we introduce the implementation of the UDA model as shown in Figure 3, which is built on a sequential variational inference module with an autoregressive hidden Markov Module for latent environment estimation. Moreover, we devise modular prior networks to estimate the prior of stationary and nonstationary latent variables.
Figure 3: The framework of the UDA model. The latent variable encoder is used to extract ${\mathbf{z}}^{s}_{1:t}$ and ${\mathbf{z}}^{e}_{1:t}$ from ${\mathbf{x}}_{1:t}$. The latent forecasting module is used to estimate ${\mathbf{z}}^{s}_{t+1:T}$ and ${\mathbf{z}}^{e}_{t+1:T}$ from ${\mathbf{z}}^{s}_{1:t}$ and ${\mathbf{z}}^{e}_{1:t}$. The future forecasting module is used for future prediction $\{\hat{x}_{t+1},\cdots,\hat{x}_{T}\}$. The historical latent environments $\{\hat{{\mathbf{e}}}_{1},\cdots,\hat{{\mathbf{e}}}_{t}\}$ are generated by the environment estimation module, and the future latent environments $\{\hat{{\mathbf{e}}}_{t+1},\cdots,\hat{{\mathbf{e}}}_{T}\}$ are generated by the environment prediction module. The nonstationary prior and the stationary prior are used to estimate the prior distribution of stationary and nonstationary latent variables for KL divergence.
@@ 4.1 Sequential Variational Inference Module for Time Series Data Modeling
Based on the data generation process in Figure 2, we first derive the evidence lower bound (ELBO) in Equation (7). Please refer to Appendix D for more details of the derivation.
@@
$\small\begin{split}\mathcal{L}_{KLD}^{s}\!&=D_{KL}\!(q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{s}_{1:t}))\\ &+\mathbb{E}_{q({\mathbf{z}}^{s}_{1:t}|{\mathbf{x}}_{1:t})}\!\Big{[}\!D_{KL}\!(q({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t})||p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}))\Big{]}\\ \mathcal{L}_{KLD}^{e}\!&=D_{KL}\!(q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})||p({\mathbf{z}}^{e}_{1:t}))\\ &+\mathbb{E}_{q({\mathbf{z}}^{e}_{1:t}|{\mathbf{x}}_{1:t})}\!\Big{[}\!D_{KL}\!(q({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})||p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t}))\Big{]},\end{split}$ (9)
in which $q({\mathbf{z}}_{1:t}^{s}|{\mathbf{x}}_{1:t}),\,\,q({\mathbf{z}}_{t+1:T}^{s}|{\mathbf{z}}_{1:t}^{s}),\,\,q({\mathbf{z}}_{1:t}^{e}|{\mathbf{x}}_{1:t})$ and $q({\mathbf{z}}_{t+1:T}^{e}|{\mathbf{z}}_{1:t}^{e})$ are used to approximate the distribution. Therefore, the aforementioned approximate functions, the historical decoder, and the future forecasting module can be formalized as follows:
$\small\begin{split}\hat{\mathbf{z}}^{e}_{1:t}&=\psi_{e}({\mathbf{x}}_{1:t};\theta_{\psi_{e}}),\quad\quad\,\,\,\,\quad\hat{\mathbf{z}}^{s}_{1:t}=\psi_{s}({\mathbf{x}}_{1:t};\theta_{\psi_{s}}),\,\,\,\,\quad\\ \hat{\mathbf{z}}^{e}_{t+1:T}&=T_{e}(\hat{\mathbf{z}}^{e}_{1:t};\theta_{T_{e}}),\,\,\,\,\quad\quad\hat{\mathbf{z}}^{s}_{t+1:T}=T_{s}(\hat{\mathbf{z}}^{s}_{1:t};\theta_{T_{s}}),\,\,\,\,\quad\\ \hat{{\mathbf{x}}}_{1:t}&=F_{x}(\hat{\mathbf{z}}^{e}_{1:t},\hat{\mathbf{z}}^{s}_{1:t};\theta_{x}),\quad\hat{{\mathbf{x}}}_{t+1:T}=F_{y}(\hat{\mathbf{z}}^{e}_{t+1:T},\hat{\mathbf{z}}^{s}_{t+1:T};\theta_{y}),\end{split}$ (10)
@@
$p({\mathbf{z}}^{s}_{1:t}),\,\,p({\mathbf{z}}^{s}_{t+1:T}|{\mathbf{z}}^{s}_{1:t}),\,\,p({\mathbf{z}}^{e}_{1:t})$ and $p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})$ are the estimated prior distribution of stationary and nonstationary latent variables, which will be introduced in the Subsection 4.2. Note that the environment estimation module does not appear in the ELBO explicitly, which is a part of the $p({\mathbf{z}}^{e}_{1:t}),p({\mathbf{z}}^{e}_{t+1:T}|{\mathbf{z}}^{e}_{1:t})$ and will be also introduced in subsection 4.2. Please refer to the Appendix H for more details of the implementation of the proposed UDA model.
4.2 Stationary and Nonstationary Priors Estimation
Previous time-series modeling methods based on the causality-based data generation processes usually require autoregressive inference and Gaussian prior (Fabius and Van Amersfoort,, 2014; Kim et al.,, 2019; Zhu et al.,, 2020). However, simply assuming the Gaussian distribution might result in suboptimal performance of disentanglement. To solve this problem, we employ the modular neural architecture to estimate the prior distribution of stationary and nonstationary latent variables.
Modular Architecture for the Stationary Prior Estimation. we first let $\{r_{i}^{s}\}$ be a set of learned inverse transition functions that take the estimated stationary latent variables and output the noise term, i.e., $\hat{\epsilon}_{t,i}^{s}=r_{i}^{s}(\hat{z}_{t,i}^{s},\hat{{\mathbf{z}}}^{s}_{t-1})$  [^2] We use the superscript symbol to denote estimated variables. and each $r_{i}^{s}$ is modeled with MLPs. Then we devise a transformation $\phi^{s}:=\{\hat{{\mathbf{z}}}^{s}_{t-1},\hat{{\mathbf{z}}}^{s}_{t}\}\rightarrow\{\hat{{\mathbf{z}}}^{s}_{t-1},\hat{\epsilon}^{s}_{t}\}$, and its Jacobian is $\small{\mathbf{J}_{\phi^{s}}=\begin{pmatrix}\mathbb{I}&0\\ *&\text{diag}\left(\frac{\partial r^{s}_{i}}{\partial\hat{z}_{t,i}^{s}}\right)\end{pmatrix}}$, where * denotes a matrix. By applying the change of variables formula, we have: